Journalist Report – March 6th

Sol 22 – “The last week’s a smooth ride”, they said!

“The Martian mountains lay all around, millions of years old. Starlight glittered on the spires of a little Martian town, no bigger than a game of chess, in the blue hills.”

– Chapter 22 of The Martian Chronicles by Ray Bradbury

I woke up this morning as my crewmates silently and sleepily stepped out of their rooms to take their daily tests and measurements, as I lay on a reclined mattress under the main Hab window. It feels strange not being able to have the same habits as the rest of my crew; as Corentin led them all through the daily workout session, I took advantage of my not being able to participate by taking my role of crew photographer close at heart.

There was no time to waste after breakfast; after being delayed twice, the third and last EVA at Candor Chasma dedicated to the photogrammetry experiment would finally take place this morning! Corentin having replaced me as the third member of the EVA crew, I was assigned the role of HabCom to monitor the EVA from the Hab. It was a quiet morning: after handling the data retrieved from the atmospheric instruments, which the researchers from CNRS are very pleased with, Alexandre played chess with Quentin to the sweet churning sound of the bread-making machine. Meanwhile, Jérémy tested all of AI4U’s functionalities. The objective is to have all crewmembers interact with the AI in order to give detailed feedback to Spoon and CNES, who have created the prototype.

After four hours of exploration, our EVA crew came back home, sweaty, exhausted, but thrilled to have finally collected all the necessary data for the photogrammetry experiment. We enjoyed a long lunch and some off-worldly conversations. It’s amazing to see how close we have become over the past few weeks, despite our differences: every single crewmember is doing their best for this last week to go as smoothly as possible.
In the afternoon, we struggled to keep to a tight schedule. Crewmembers are taking turns using AI4U, and getting their cognitive tests and physiological examinations done. Alice has started comparing the performances of the photogrammetry teams, who are feeling quite competitive…

Tomorrow morning, I’ll be travelling back to Earth with Alice and Corentin, and possibly getting surgery. Jérémy will be writing the reports for the next few Sols, and I can only hope to be back before the mission ends.

In any case, this evening, although bittersweet, will most certainly reflect the team’s spirit during the entire mission: quick with a joke, never sparing with kind words.

EVA Report – March 6th

Crew 275 EVA Report 06Mar2023

EVA # 18

Author: Adrien Tison, EVA Leader, and Crew Botanist

Purpose of EVA: The purpose of the EVA was to perform the photogrammetry experiments with the 3D map.

First, as it was on the way to Candor Chasma, we changed the batteries and re-installed the LOAC.

We then went to Candor Chasma, where the subject of the experiment tried to find some specific checkpoints, while the operators of the experiment onsite took time measurements and drew their trajectory on a map. In addition, the subject was equipped with a step counter. Indeed, as the Crew Journalist is injured, we performed the experiment with only one subject.

Start time: 9:05

End time: 12:39


9:05 – Airlock closed, depressurization starting

9:10 – Exiting Airlock, airlock closed

9:13 – Departure from Hab

9:16 – Arrival at the instruments’ site

9:25 – Batteries are being changed and data retrieved. MegaAres was off when we arrived

9:37 – Leaving the site

9:39 – Departure with the rovers

9:43 – Arrived at North Ridge

9:44 – Leaving the rovers and walking

9:54 – Radio check from HabCom– OK – HabCom is received very well on EVA

Leader’s side – Still on their way to Candor Chasma

10:05 – Brief contact with EVA leader, switching to channel 2

10:16 – Photogrammetry experiment starting, EVA leader has to search for the checkpoints

11:14 – Experiment over, exploration of Candor Chasma

11:51 – EVA Crew heading back to the rovers

12:32 – EVA Crew back at the rovers

12:36 – Passing Marble Ritual

12:38 – Rovers parked at the Hab

12:39 – Crew in Airlock

To conclude briefly, everything went to plan, batteries have been changed, LOAC is now in place and data have been retrieved. The photogrammetry experiment went perfectly well, all checkpoints were found quite quickly.

Destination: First we will stop where the MegaAres and the METEO station are

located; we will then go to Candor Chasma.

Coordinates: N4251000 E521500

EVA Participants: EVA Leader: Adrien Tison (Crew Botanist & Subject), Alice Chapiron (Crew Scientist & Operator), and Corentin Senaux (HSO & Operator)

HAB COM : Marie Delaroche (Crew Journalist)

Road(s) and routes per MDRS Map – Cow Dung Road until the intersection with “Gateway to Candor”, we then walked eastwards until Candor Chasma for an estimated 2.5 km from the rovers.

Mode of travel – Driving and walking

Vehicles we used – Opportunity and Spirit

GreenHab Report – March 6th

GreenHab Officer: Adrien Tison

Environmental control: Heater

Average Temperature: 75.4°F

Average humidity: 23%

Hours of supplemental light: 0 hour

Daily water usage for crops: 12.68 gallons

Daily water usage for research and/or other purposes: 0 gallons

Water in the blue tank: 86.67 gallons

Time(s) of watering for crops: 9:45 & 15:45

Changes to crops: N/A

Narrative: Today I was a member of the morning EVA, thus, it was the crew astronomer who watered plants and crops in the morning.

In the afternoon, I went back to water them again and saw tomatoes ready to be harvested. I then harvested them after watering everything.

I also ran some water tests on my aquaponics.

Harvest: 808 g of tomatoes

Support/supplies needed: N/A

Operations Report – March 6th

Crew 275 Operations Report 06Mar2023
SOL: 22

Name of person filing report: Quentin Royer
Non-nominal systems:
• Toilet
• Spacesuit n°9
• Robotic observatory
• Spacesuit n°4 and a helmet

Notes on non-nominal systems:
• Spacesuit n°4: When the helmet of the spacesuit was removed, as a Crew Member unscrewed the plastic tubes linking the helmet and the suit, the screw remained blocked, and the two parts are very firmly screwed (it was probably initially badly screwed). Therefore, we had to remove the plastic piece from the spacesuit, and it is now impossible to separate it from the plastic part attached to the helmet.

Spirit rover used: Yes
• Hours: 209.9
• Beginning charge: 100%
• Ending charge: 73%
• Currently charging: Yes

Opportunity rover used: Yes
• Hours: 113.9
• Beginning charge: 100%
• Ending charge: 81%
• Currently charging: Yes

Curiosity rover used: No
• Hours: 221.7
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Perseverance rover used: No
• Hours: 255.9
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

General notes on rovers: Nothing to report.

Summary of Hab operations: The Crew conducted routine operations.

• WATER USE: 12.87
• Water: 365.65 gallons
• Static tank pipe heater: On
• Static tank heater: On
• Toilet tank emptied: Yes

Summary of internet: Starlink (MDRSCampus-Guest) is used during COMMS and during the day for emails and to handle the medical emergency. MDRScampus is used for experimental purpose.

Summary of suits and radios: Nothing to report.

Summary of GreenHab operations: The Crew botanist conducted routine operations.
• WATER USE: 12.68 gallons
• Heater: On
• Supplemental light: No
• Harvest: 808 g of tomatoes

Summary of ScienceDome operations: No operations were conducted in the ScienceDome.
• Dual split: Heat / From 10 pm to 7 am

Summary of RAM operations: No operations were conducted in the RAM today.

Summary of any observatory issues: The robotic observatory is non-nominal.

Summary of health and safety issues: None.

Questions, concerns, and requests to Mission Support: None.

Sol Summary Report – March 6th

Crew 275 Sol Summary Report 06Mar2023

Sol: 22

Summary Title: “The last week’s a smooth ride”, they said!

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

Another 4-hour EVA left three crewmembers very tired today. They went to Candor Chasma and successfully performed the photogrammetry EVA, based on a 3D render of the region. Only one crewmember could take part as a subject, since the other subject was injured. Before reaching Candor Chasma, they stopped to install the LOAC, and then to change the batteries and retrieve the data of the other atmospheric instruments. They noticed that MegaAres was off, as it ran out of battery. We have noticed that about 3 hours of recording are missing. This is due to the fact that we did not have enough time and enough chargers to completely charge the batteries of all the atmospheric instruments.

With the GreenHab officer in EVA, the Crew Astronomer watered the crops. We then spent some additional time in the GreenHab to perform a few cognitive tests, while the temperature there was very close to 100°F/40°C.

The injured crewmember is feeling slightly better, and we managed to get an appointment with a surgeon for her. We will drive her to Richfield tomorrow. However, the fact that a few crewmembers will not be at the MDRS required us to modify the schedule and to perform additional activities today.

Look Ahead Plan:

Today, we managed to get an appointment with a surgeon in Richfield. The injured crewmembers will have to be at the Sevier Valley hospital tomorrow (Sol 23) at 10:40 for a second X-ray and then to see the surgeon. We requested Mission Support to use the CrewCar to do so. The plan is to leave between 7:00 and 7:30 with the Crew Journalist, the Crew Scientist, and the Health & Safety Office, so that two drivers can accompany the injured crewmember. In the absence of the Health & Safety Officer, the Commander will take this role. The tasks of the Crew Scientist and of the Crew Journalist will be shared between the four crewmembers remaining at the MDRS.

A media visit is planned for tomorrow (Sol 23) with a crew from France Televisions. They will arrive at 9:00 and leave at 17:00. In the morning, we plan to have them following us during our EVA, first to change the battery and retrieve the data from the atmospheric instruments, and then to shoot nice videos at North Ridge. In the afternoon, we will show them the different buildings and the rest of our experiments.

We would also like to perform an exploration EVA on Thursday (Sol 25) and another EVA dedicated to the disassembling of the atmospheric instruments on Friday (Sol 26).

Anomalies in work: None.

Weather: Windy, low 1°C/34°F, high 12°C/54°F.

Crew Physical Status: One crewmember injured (broken collarbone / clavicula). Appointment with surgeon is planned tomorrow (Sol 23) at 11:10 in Richfield. Surgery should soon follow.

EVA: EVA # 18 dedicated to the photogrammetry experiment based on a 3D render was successfully performed. Some anomalies were observed regarding the batteries of the atmospheric instruments.

Reports to be filed: Sol Summary, EVA Report, EVA Request, Operations Report, Journalist Report, GreenHab Report, HSO Report.

Support Requested: None.

Journalist Report – March 5th

Sol 21 – Sunday Slumber

“They all came out and looked at the sky that night. They left their suppers or their washing up or their dressing for the show and they came out upon their now-not-quite-as-new porches and watched the green star of Earth there.”

– – Chapter 21 of The Martian Chronicles by Ray Bradbury

Our third Sunday in the MDRS started out as usual: with a little more sleep for the entire crew. After yesterday’s emotional rollercoaster, somewhat rested, we tucked into some more pancakes prepared by Alice. We watched and laughed as she tested AI4U’s joke functionalities; over the past week, Alexandre has continued running tests on the program to help the CNES researchers make the right changes to it, with future space missions in mind.

Our third Sunday in the MDRS continued not exactly as usual: two crewmembers were scheduled to go on an EVA. As Sol 20’s outing was cancelled, the batteries on all atmospheric instruments had not been changed and had been left in the cold; Alexandre had therefore requested a Sunday EVA. Despite the 20 mph winds, which the crew is now used to hear blowing around the station, Alice and Alexandre managed to change all three batteries and retrieve the data. Unfortunately, the LOAC had to be brought home once again to be fixed, as it has not been collecting any data for the past Sol…

With everyone back at the Hab for the afternoon, everything started to feel like a real Sunday at MDRS, and we were glad to have more time to rest after Friday night’s events. Naps were taken, movies were watched, books were read.

Despite not being able to work on his astronomy project due to a malfunctioning robotic observatory, Alexandre has taken up astrophotography. Along with Adrien, he set up a tripod and camera under the glass dome of the observatory to take advantage of the beautiful Martian skies. We have not yet been able to see Earth from up here, we are still actively looking…

EVA Report – March 5th

Crew 275 EVA Report 05-03-2023

EVA # 17

Author: Alexandre VINAS (crew astronomer and EVA leader)

Purpose of EVA: Change of battery for all the instruments, as well as data retrieval.

Start time: 10:18

End time: 11:30


10:18 Airlock closed

10:28 EVA crew arrived at the destination.

MegaAres instrument was off. Its 12V battery had a level of 8.6V.

LOAC malfunctionned. Its battery was fully charged but the instrument was off (even if it was turned on during the last EVA). LOAC was removed.

11:02 Change of the batteries of PurpleAir and data retrieval.

11:18 Change of the batteries of Field mill and data retrieval.

11:25 EVA crew left the site of the instruments.

11:30 EVA crew back at the rovers.

11:45 EVA crew back in the hab.

Destination: Between the Hab and Marble Ritual near the road.

Coordinates (use UTM WGS 84): N4251000 E518500

Participants: EVA Leader: Alexandre VINAS (Crew astronomer), Alice Chapiron (Crew Scientist)

HAB COM: Quentin Royer (Crew Engineer)

Road(s) and routes per MDRS Map: Drive Cow Dung Road 0110 to a place just before Marble Ritual (see map on the attached picture)

Mode of travel: Curiosity

GreenHab Report – March 5th

GreenHab Officer: Adrien Tison

Environmental control: Heater

Average Temperature: 78.5°F

Average humidity: 28%

Hours of supplemental light: 0 hour

Daily water usage for crops: 9.51 gallons

Daily water usage for research and/or other purposes: 0 gallons

Water in the blue tank: 99.35 gallons

Time(s) of watering for crops: 10:45

Changes to crops: I finally harvested the last greens and then planted new seeds in this planter next to the blue tank

Narrative: I went to water and maintain crops as my daily operations.

Thankfully the dirt wasn’t too dry considering I only went once yesterday.

Apart from that, I harvested the greens located in the planter next to the blue tank and planted new seeds. That way both planters may give some good greens for next crews!

Harvest: 81g of greens

Support/supplies needed: N/A

Copyright © The Mars Society. All rights reserved. | Main Site